OpenStudy (destinyyyy):

I have a "few" word problems. Could someone help me out? Its been awhile since Ive done one.

OpenStudy (destinyyyy):

I have no examples. So please bare with me.

OpenStudy (destinyyyy):

A new car worth $24,000 is depreciating in value by $3,000 per year. Complete A- B. A. Write the formula that models the cars value, y, in dollars, after x years.

OpenStudy (destinyyyy):

This is question 5.

OpenStudy (lfreeman1285):

A. If the car's value declines by 3000/year, then this can be represented by 3000x, and because the car originally cost 24000, the equation would be y=24000-3000x

OpenStudy (destinyyyy):

Okay

OpenStudy (destinyyyy):

B. Use the formula from part A to determine after how many years the cars value will be $15,000.

OpenStudy (lfreeman1285):

Here, you would just plug in 15000 for y. 15000=24000-3000x There are a number of ways you can go about solving this, but in the end x has to be isolated on one side of the equation. Personally, I like to subtract and then divide. First, I would subtract 24000 from both sides -9000=-3000x And then I divide by -3000. 3=x So three years is your answer.

OpenStudy (destinyyyy):

Okay. Thank you. Would you mind helping with a few more?

OpenStudy (lfreeman1285):

Sure!

OpenStudy (lfreeman1285):

do you know if we should assume that a month is 30 days?

OpenStudy (destinyyyy):

I honestly don't know

OpenStudy (lfreeman1285):

@Vocaloid what do you think?

OpenStudy (destinyyyy):

The help me solve this button says " Here, x is the number of times in a month the bus is used."

OpenStudy (lfreeman1285):

Maybe its 1.75x=30-0.75x? That would make the most sense to me.

OpenStudy (destinyyyy):

Okay.. I have new numbers now since I clicked help me solve this... 1.50x= 24x- .50x

OpenStudy (destinyyyy):

= 24 not 24x

OpenStudy (destinyyyy):

I got x=12